# Expected Coordinate Improvement for High-Dimensional Bayesian Optimization

### v1

@@ 4.1 Experiment Settings
From the table we can see that our proposed ECI-BO obtains significantly better results than the standard BO on most of the test problems. For example, on the unimodal problem $f_{1}$, the average result found by the standard BO is 1.16E+10 while the average result found by the proposed ECI-BO is 2.76E+07. On the multimodal problem $f_{10}$ the numbers are 3.47E+04 and 1.75E+04, on the hybrid problem $f_{20}$ the numbers are 8.85E+03 and 6.04E+03, and on the composition problem $f_{30}$ the numbers are 7.95E+09 and 1.19E+09. Overall, our proposed ECI-BO finds better solutions on twenty-four test problems, worse solutions on two test problems and similar solutions on three test problems. The major difference between these two algorithms is that the proposed ECI-BO searches along one coordinate to find an infill solution in each iteration while the standard BO searches in the original design space for finding an infill solution. By iterating over the coordinates, our proposed ECI-BO is able to improve the solution gradually. At the end of 1000 evaluations, our proposed ECI-BO is able to find better solutions on most of the test problems. The experiment results can empirically prove the effectiveness of the proposed expected coordinate improvement approach.
The iteration histories of the two compared algorithms on six selected functions $f_{1}$, $f_{6}$, $f_{12}$, $f_{18}$, $f_{24}$ and $f_{30}$ are plotted in Fig. 5, where the horizontal axis is the number of function evaluations and the vertical axis is the current minimum objective value. The median, first quartile and the third quartile of the 30 runs are shown in the figures. From the iteration histories we can find a very interesting phenomenon. The standard BO converges faster than the proposed ECI-BO at the beginning of iterations, but converges slower than the proposed ECI-BO at the end of the iterations on most of the test problems. The standard BO searches in the original space to find new solutions while the proposed ECI-BO searches along one coordinate. Therefore, the standard BO is able to find better solutions than the ECI-BO at the beginning of the iterations. However, as the iteration goes on, most of the promising areas have been explored and exploitation is needed for further improvement. The standard BO still searches in the whole design space to find promising solutions, which often brings very little improvement as the iteration continues. On the contrary, the proposed ECI-BO searches along one coordinate at a time to find new solutions. Although it converges slower than the standard BO at the beginning of iterations when exploration is needed, it is able to continually find good solutions at the middle and at end of iterations when exploitation is urgently needed. At the end of iterations, the proposed ECI-BO often finds better solutions than the standard BO approach. Since the standard BO has faster convergence speed at the beginning of iterations, it is still preferred when the computation budget is extremely limited. But when the computation budget is fairly enough, the proposed ECI-BO should be preferred.
At last, we compare the computational time of the two algorithms. The computational time of them within one iteration consists of the time for training the GP model, the time for optimizing the acquisition function, and the time for evaluating the new point. The major difference between them is in optimizing the acquisition function. The proposed ECI-BO maximizes the 1-D ECI function to query a new point in each iteration while the standard BO maximizes the $d$-dimensional EI function to query a new point. As a result, the infill optimization problem of the proposed ECI-BO is often faster to solve than that of the standard BO algorithm. The running time of the two compared algorithms is plotted in Fig. 6, where the average computation time of 30 runs is shown. We can see that our proposed ECI-BO is significantly faster than the standard BO on 100-D problems. In summary, the proposed expected coordinate improvement approach is able to improve the standard Bayesian optimization algorithm’s optimization efficiency and reduce its computational cost at the same time.
Figure 6: Running time of the proposed ECI-BO and the standard BO.
4.2 Comparison with High-dimensional BOs
To verify the effectiveness of the proposed approach, we compare our ECI-BO wit five high-dimensional BOs. The compared algorithms are described in the following.
  1. 1.
    Add-GP-UCB: The Add-GP-UCB algorithm [33] is the most popular high-dimensional BOs with the additive structural assumption. It decomposes the original problem into multiple low-dimensional sub-problems, solves these sub-problems and combines the solutions of the sub-problems into a high-dimensional solution. The low-dimensional space is also set to five in the experiments. The open source code of Add-GP-UCB  [^1] https://github.com/kirthevasank/add-gp-banditss is used in the comparison.
  2. 2.
    Dropout: The Dropout approach [18] randomly selects a subset of variables to perform BO, and combine the solutions of the sub-problem with existing solutions to form a high-dimensional solution. Three different strategies for forming the high-dimensional solution are proposed in the work [18], and the copy strategy is used in the comparison. The dimension of the subspace is also set to five in the experiments. The open source code of the Dropout approach  [^2] https://github.com/licheng0794/highdimBO_dropout is used in the comparison.
  3. 3.
    CoordinateLineBO: The CoordinateLineBO approach [39] randomly selects a coordinate to optimize in each iteration. The coordinates of the infill solution are the same as the current best solution except for the optimized coordinate. We implement the CoordinateLineBO algorithm in Matlab according to the work [39]. We use the genetic algorithm with population size of 10 and maximal generation of 20 to optimize the selected coordinate in each iteration for the CoordinateLineBO approach.
  4. 4.
    TuRBO: The TuRBO [37] uses the trust region approach to improve local search ability of Bayesian optimization algorithm on high-dimensional problems. The Thompson sampling is employed in TuRBO to select a batch of candidates to evaluate in each iteration. The number of trust regions is set to five and the number of batch evaluations is set to ten in the experiment. The open source code of TuRBO  [^3] https://github.com/uber-research/TuRBO is used in the comparison.
  5. 5.
    MCTS-VS: The MCTS-VS approach [19] is a high-dimensional BO with variable selection strategy. The Monte Carlo tree search method in employed iteratively to select a subset of variables to optimize in each iteration. The open source code of MCTS-VS  [^4] https://github.com/lamda-bbo/MCTS-VS is used in the comparison.
The experiment results on the 100-D problems are shown in Table 2, where the average and standard deviation values of 30 runs are given. We also conduct Wilcoxon signed rank test, and use $+$, -- and $\approx$ to donate that our ECI-BO achieves significantly better, significantly worse and similar results compared with the competing algorithms.
Table 2: Average and standard derivation results of the compared high-dimensional BOs
$f$ Add-GP-UCB Dropout Coordinate LineBO TuRBO MCTR-VS ECI-BO
$f_{1}$ 2.35E+11 $+$ 9.60E+10 $+$ 4.20E+09 $+$ 8.32E+10 $+$ 1.39E+11 $+$ 2.76E+07
$f_{3}$ 4.45E+09 $\approx$ 9.54E+05 $\approx$ 9.63E+05 $\approx$ 9.54E+05 $\approx$ 2.36E+06 $+$ 9.70E+05
$f_{4}$ 7.09E+04 $+$ 1.67E+04 $+$ 6.81E+03 $+$ 3.24E+04 $+$ 2.51E+04 $+$ 3.37E+03
$f_{5}$ 2.21E+03 $+$ 2.02E+03 $+$ 1.69E+03 $+$ 1.74E+03 $+$ 1.67E+03 $+$ 1.42E+03
$f_{6}$ 7.24E+02 $+$ 6.72E+02 $+$ 6.42E+02 $+$ 6.55E+02 $+$ 6.60E+02 $+$ 6.12E+02
$f_{7}$ 4.49E+03 $+$ 4.15E+03 $+$ 2.35E+03 $+$ 3.60E+03 $+$ 4.73E+03 $+$ 1.81E+03
$f_{8}$ 2.58E+03 $+$ 2.37E+03 $+$ 2.05E+03 $+$ 2.05E+03 $+$ 1.96E+03 $+$ 1.71E+03
$f_{9}$ 9.13E+04 $+$ 1.05E+05 $+$ 8.26E+04 $+$ 8.56E+04 $+$ 5.02E+04 $\approx$ 5.68E+04
$f_{10}$ 3.29E+04 $+$ 2.60E+04 $+$ 2.30E+04 $+$ 2.84E+04 $+$ 3.42E+04 $+$ 1.75E+04
$f_{11}$ 7.04E+06 $+$ 1.83E+05 -- 2.89E+05 -- 3.36E+05 $\approx$ 4.52E+05 $+$ 4.46E+05
$f_{12}$ 2.14E+11 $+$ 2.36E+10 $+$ 4.05E+09 $+$ 2.13E+10 $+$ 5.25E+10 $+$ 7.49E+08
$f_{13}$ 4.46E+10 $+$ 3.83E+09 $+$ 1.68E+09 $+$ 2.28E+09 $+$ 1.10E+10 $+$ 2.30E+08
$f_{14}$ 3.49E+08 $+$ 5.83E+07 -- 9.01E+07 $+$ 7.67E+07 $\approx$ 1.30E+08 $+$ 7.20E+07
$f_{15}$ 3.03E+10 $+$ 9.14E+08 $+$ 1.60E+09 $+$ 5.51E+08 $\approx$ 5.06E+09 $+$ 5.38E+08
$f_{16}$ 1.53E+04 $+$ 9.70E+03 -- 1.14E+04 $+$ 1.42E+04 $+$ 1.27E+04 $+$ 1.05E+04
$f_{17}$ 3.98E+07 $+$ 1.58E+04 -- 1.89E+06 $\approx$ 1.01E+05 -- 3.60E+04 -- 2.07E+06
$f_{18}$ 6.86E+08 $+$ 5.65E+07 -- 1.08E+08 $\approx$ 1.10E+08 $\approx$ 2.61E+08 $+$ 1.06E+08
$f_{19}$ 2.93E+10 $+$ 1.10E+09 $+$ 9.95E+08 $+$ 5.16E+08 $+$ 5.78E+09 $+$ 1.28E+08
$f_{20}$ 8.63E+03 $+$ 6.33E+03 $\approx$ 7.09E+03 $+$ 5.84E+03 $\approx$ 8.45E+03 $+$ 6.04E+03
$f_{21}$ 4.19E+03 $+$ 3.87E+03 $+$ 3.56E+03 $+$ 3.90E+03 $+$ 3.59E+03 $+$ 3.40E+03
$f_{22}$ 3.60E+04 $+$ 2.91E+04 $+$ 2.44E+04 $+$ 3.06E+04 $+$ 3.64E+04 $+$ 1.96E+04
$f_{23}$ 5.06E+03 $+$ 4.48E+03 $+$ 3.93E+03 $+$ 4.48E+03 $+$ 4.22E+03 $+$ 3.74E+03
$f_{24}$ 6.38E+03 $+$ 5.95E+03 $+$ 4.50E+03 $+$ 5.23E+03 $+$ 5.26E+03 $+$ 4.29E+03
$f_{25}$ 2.40E+04 $+$ 1.27E+04 $+$ 7.08E+03 $+$ 2.06E+04 $+$ 1.99E+04 $+$ 4.98E+03
$f_{26}$ 3.86E+04 $+$ 2.92E+04 $+$ 1.95E+04 $+$ 2.94E+04 $+$ 2.40E+04 $+$ 1.81E+04
$f_{27}$ 7.14E+03 $+$ 5.35E+03 $+$ 4.01E+03 $+$ 5.17E+03 $+$ 5.19E+03 $+$ 3.81E+03
$f_{28}$ 2.38E+04 $+$ 1.67E+04 $+$ 1.50E+04 $+$ 1.70E+04 $+$ 2.22E+04 $+$ 8.81E+03
$f_{29}$ 6.25E+06 $+$ 1.22E+04 -- 2.18E+05 -- 3.63E+04 -- 2.76E+04 -- 4.26E+05
$f_{30}$ 4.00E+10 $+$ 2.87E+09 $+$ 2.26E+09 $+$ 1.70E+09 $+$ 7.78E+09 $+$ 1.19E+09
$+$/$\approx$/-- 28/1/0 21/2/6 24/3/2 21/6/2 26/1/2 N.A.
@@ 4.2 Comparison with High-dimensional BOs
The Add-GP-UCB [33] assumes the objective function is a summation of multiple low-dimensional functions. However, the problems in CEC 2017 test suite do not meet this assumption. In comparison, our proposed ECI-BO does not make any structural assumption about the objective function we are optimizing. From Table 2, we can find that the Add-GP-UCB can not achieve satisfying results on the test problems. Our proposed ECI-BO performs significantly better than the Add-GP-UCB on twenty-eight of the twenty-nine problems.
@@
The Dropout approach [18] randomly selects a few variables to optimize in each iteration. When evaluating the new solutions, this approach uses the values of existing points to fill the unselected variables. From the table we can find that our proposed ECI-BO performs significantly better on twenty-one problems, and significantly worse on six problems compared with the Dropout approach.
@@
The CoordinateLineBO [39] is the most similar approach to our proposed ECI-BO. Both the CoordinateLineBO and our ECI-BO select one variable to optimize in each iteration. However, there are two major differences between them. First, our proposed algorithm iterates the coordinates over permutations while the the CoordinateLineBO iterates the coordinates randomly. When we iterates $d$ times, all the variables will be covered by the proposed approach, but some variables may not be covered by the CoordindateLineBO approach since some variables might be selected multiple times. Second, the CoordinateLineBO treats the variables equally, while our proposed ECI-BO learns the effectiveness of the variables based on their ECI values. The experiment results in Table 2 show that the proposed ECI-BO performs significantly better than the CoordinateLineBO approach on twenty-four problems. This can empirically prove the effectiveness of the proposed expected coordinate improvement approach.
@@
The TuRBO [37] employees the trust region strategy in the BO framework to improve the local search ability of BO on high-dimensional problems. Compared with the TuRBO, our ECI-BO finds better solutions on twenty-one problems, similar solutions on six problems, and worse solutions on two problems. This means our proposed expected coordinate improvement approach is more effective than the trust region approach on the 100-D CEC 2017 test suite.
@@
The MCTS-VS [19] is a recently developed variable selection strategy for high-dimensional Bayesian optimization. It employees the Monte Carlo tree search method to partition the variables into important and unimportant ones, and only optimizes the important variables [19]. The experiment results show that our proposed ECI-BO performs significantly better than the MCTS-VS approach on most of the test problems. This indicates that our ECI-BO is more suitable for problems whose variables are all effective to the objective function.

### v2

@@ 4.2 Comparison with Bayesian Optimization Algorithms
First, we compare our proposed ECI-BO with the standard BO algorithm. From the tables, we can see that our proposed ECI-BO obtains significantly better results than the standard BO on most of the test problems. Overall, our proposed ECI-BO finds better solutions on eighteen of the twenty analytical test problems, fifty-nine of the eighty-four CEC 2013 problems, and seventy-three of the eighty-seven CEC 2017 test problems . The major difference between these two algorithms is that the proposed ECI-BO searches along one coordinate to find an infill solution in each iteration while the standard BO searches in the original design space for finding an infill solution. On high-dimensional problems, it is very challenging for the standard BO to find good solutions in a very large search space. At the end of 1000 evaluations, our proposed ECI-BO is able to find better solutions on most of the test problems. The experiment results can empirically prove the effectiveness of the proposed expected coordinate improvement approach.
@@
The Add-GP-UCB [38] assumes the objective function is a summation of multiple low-dimensional functions. However, most of the the selected problems do not meet this assumption. In comparison, our proposed ECI-BO does not make any structural assumption about the objective function we are optimizing. From the tables, we can find that the Add-GP-UCB can not achieve satisfying results on the test problems. Our proposed ECI-BO performs significantly better than the Add-GP-UCB on most of the benchmark problems.
@@
The Dropout approach [23] randomly selects a few variables to optimize in each iteration. When evaluating the new solutions, this approach uses the values of existing points to fill the unselected variables. The Dropout approach also trains the model in the selected subspace, which might result in poor GP models. In comparison, our approach trains the GP model in the original design space. This might be the reason why our proposed approach has better performance than the Dropout approach. From the tables we can find that our proposed ECI-BO performs significantly better on eighteen analytical problems, fifty-two CEC 2013 problems and sixty-five CEC 2017 problems compared with the Dropout approach.
@@
The CoordinateLineBO [44] is the most similar approach to our proposed ECI-BO. Both the CoordinateLineBO and our ECI-BO select one variable to optimize in each iteration. However, there are two major differences between them. First, our proposed algorithm iterates the coordinates over permutations while the the CoordinateLineBO iterates the coordinates randomly. When we iterates $d$ times, all the variables will be covered by the proposed approach, but some variables may not be covered by the CoordindateLineBO approach since some variables might be selected multiple times. Second, the CoordinateLineBO treats the variables equally, while our proposed ECI-BO learns the effectiveness of the variables based on their ECI values. The experiment results in the tables show that the proposed ECI-BO performs significantly better than the CoordinateLineBO approach on most of the test problems. This can empirically prove the effectiveness of the proposed expected coordinate improvement approach.
@@
The TuRBO [42] employees the trust region strategy in the BO framework to improve the local search ability of BO on high-dimensional problems. Although TuRBO reduces the search space to local regions, the local regions are still high-dimensional spaces. Searching in the high-dimensional local spaces is still challenging. Compared with the TuRBO, our ECI-BO finds better solutions on the majority of the test problems. This means our proposed expected coordinate improvement approach is more effective than the trust region approach on the test problems.
@@
The MCTS-VS [24] is a recently developed variable selection strategy for high-dimensional optimization. It employees the Monte Carlo tree search method to partition the variables into important and unimportant ones, and only optimizes the important variables [24]. For problems whose variables are all important, this approach might not be very efficient since most of the variables are not optimized after the variable selection. The experiment results show that our proposed ECI-BO performs significantly better than the MCTS-VS approach on most of the problems. This indicates that our ECI-BO is more suitable for problems whose variables are all effective to the objective function.
Refer to caption (a) 30-D Ellipsoid Refer to caption (b) 50-D Ellipsoid Refer to caption (c) 100-D Ellipsoid Refer to caption (d) 200-D Ellipsoid Refer to caption (e) 30-D Rosenbrock Refer to caption (f) 50-D Rosenbrock Refer to caption (g) 100-D Rosenbrock Refer to caption (h) 200-D Rosenbrock Refer to caption (i) 30-D Ackley Refer to caption (j) 50-D Ackley Refer to caption (k) 100-D Ackley Refer to caption (l) 200-D Ackley Refer to caption (m) 30-D Griewank Refer to caption (n) 50-D Griewank Refer to caption (o) 100-D Griewank Refer to caption (p) 200-D Griewank Refer to caption (q) 30-D Rastrigin Refer to caption (r) 50-D Rastrigin Refer to caption (s) 100-D Rastrigin Refer to caption (t) 200-D Rastrigin